OpenStudy (anonymous):

5. The sum of the digits of a two- digit number is 12. When the digits are reversed, the new number obtained is 15 more than twice the original number. Find the original number.

terenzreignz (terenzreignz):

Dare I even start, or do you already have your answer? XD

OpenStudy (anonymous):

i'm starting.. but i badly need a helping hand :(

OpenStudy (anonymous):

what numbers can equal 12?

OpenStudy (anonymous):

single digit numbers*

OpenStudy (anonymous):

3+9

terenzreignz (terenzreignz):

Let's let x and y be the digits of your original number such that x is the tens digit. then x + y = 12, yes?

OpenStudy (anonymous):

cant be 39, 93 is more than 15 more

OpenStudy (anonymous):

5+7?

terenzreignz (terenzreignz):

Don't guess. We'll do this systematically... we know x + y = 12, now what other expression can we derive about x and y?

terenzreignz (terenzreignz):

As for the numbers themselves, if x is the tens digit, then the first number is 10x + y, correct?

OpenStudy (anonymous):

ya..

terenzreignz (terenzreignz):

Now, if we reverse the digits, then the new number is 10y + x, yes?

terenzreignz (terenzreignz):

Now this new number is \(\color{red}{\text{15 more than TWICE the original number}}\)

terenzreignz (terenzreignz):

So, in maths terms this is \[\Large 10y+x = 2[10x+y]+15\]

OpenStudy (anonymous):

then we proceed... 10y + x = 20x + 2y + 15

terenzreignz (terenzreignz):

Yes. Simplify that, and couple it with x+y = 12 to form a system, and then just solve it.

OpenStudy (anonymous):

transposition will do?

terenzreignz (terenzreignz):

Whatever method appeals to you :P

terenzreignz (terenzreignz):

If you like, you can directly replace the x in that longer equation with 12-y, and then just solve for y.

OpenStudy (anonymous):

10y + x = 20x + 2y + 15 substitute x x=12-y 10y - (12 - y) = 20(12 - y) + 2y + 15 10y + y - 12 =240 + 20y +15 11y - 12 = 20y + 255 11y - 20y = 255+12 -9y = 267 y = -29.666666666 would be badly wrong or i'm just making it worst..

terenzreignz (terenzreignz):

There's something amiss... hang on.

OpenStudy (anonymous):

just consider the original number like xy where x is number on tens place and y on units place...now apply all the conditions on x and y and solve for them

terenzreignz (terenzreignz):

A lot of mistakes, and they all stem from these two lines. 10y - (12 - y) = 20(12 - y) + 2y + 15 10y + y - 12 =240 + 20y +15 First of all, it was 10y + x = 20x + 2y + 15 So after replacing x with 12-y you get 10y + (12 - y) + 20(12 - y) + 2y + 15 Why don't you work from here^ instead?

OpenStudy (anonymous):

is it really 10y + (12 - y) + 20(12 - y) + 2y + 15 or it should be 10y + (12 - y) = 20(12 - y) + 2y + 15..

terenzreignz (terenzreignz):

sorry, typo. Proceed

OpenStudy (anonymous):

10y + 12 - y = 240 - 20y + 2y + 15 9y + 12 = 255 - 18y 9y + 18y = 255 - 12 27y = 243 y = 9 x = 12 - y x = 12 - 9 x = 3 the number is 39 2(39) + 15 = 93 there thanks a lot master :D

terenzreignz (terenzreignz):

Much better.